OpenStudy (anonymous):

Probability: exponential random variable. Profit on a computer is $350. Suppose a warranty for one year, and the time between two failures of a computer is exponential with mean 18 months. If it costs $40 to repair a failed computer, what is the expected profit per computer? (Set up to follow)

OpenStudy (anonymous):

Let X be an exponential random variable denoting the time until failure of a computer. E[X] = 18 months, therefore λ = 1/E[X] = 1/18 Warranty repairs must be made when X < 12 P (X<12) = 1-e-λ*12 = 1 - 1/e^(2⁄3) ≈ .48658 So, I think it should be Profit = $350 - $40 (Probability of warranty), but others have said that it's just Profit = $350 - $40*2/3; where 2/3 is the expected failure. Can anyone help clarify for me?